OpenStudy (anonymous):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Pies are ordered from the Pizza Palace according to their diameter. Large:16 inches Medium: 14inches Small: 12inches How many more square inches are in the large pizza than in the small pizza? 12.56 sq in 47.10 sq in 87.92 sq in 351.68 sq in

OpenStudy (anonymous):

please help me dont go

OpenStudy (anonymous):

to find area use \[\pi r^2\] where r is the radius. so larger pizza has area \[\pi\times 8^2=64\pi\] smaller has area \[pi\times 6^2=36\pi\]

OpenStudy (anonymous):

use a calculator to subtract \[64\pi-36\pi=28\pi=87.96...\]

OpenStudy (anonymous):

please help me dont go

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please help me dont go

OpenStudy (anonymous):

so how do i find the answer

OpenStudy (anonymous):

you need a calculator if you want to make this is a decimal

OpenStudy (anonymous):

i got 87.96

OpenStudy (anonymous):

well, thanks i will try it thanks. give me a medal please i gave u one too

OpenStudy (anonymous):

Dale's Car Rentals charges a flat rate of $25 per day plus $2.50 per mile. If the number of miles is represented by "m", which algebraic expression represents the overall rate for the day using "m"? Answer 27.50m 25 + 2.50m 25m + 2.50 22.50m

OpenStudy (anonymous):

what are you doing?

OpenStudy (anonymous):

@smartbutlazy what are you doing in my question?

OpenStudy (anonymous):

Dale decides to take his 7 year old son to a ballgame. The price for the adult ticket is 15 dollars more than twice the price of the child's ticket. If you represent the price for the child's ticket using the variable "x", how would you write the algebraic expression for the adult ticket price? Answer 2x + 15 2x + 30 30x x + 30